# Algorithmic Collusion And The Minimum Price Markov Game

### v1

@@ Abstract
This paper introduces the Minimum Price Markov Game (MPMG), a dynamic variant of the Prisoner’s Dilemma. The MPMG serves as a theoretical model and reasonable approximation of real-world first-price sealed-bid public auctions that follow the minimum price rule. The goal is to provide researchers and practitioners with a framework to study market fairness and regulation in both digitized and non-digitized public procurement processes, amidst growing concerns about algorithmic collusion in online markets. We demonstrate, using multi-agent reinforcement learning-driven artificial agents, that algorithmic tacit coordination is difficult to achieve in the MPMG when cooperation is not explicitly engineered. Paradoxically, our results highlight the robustness of the minimum price rule in an auction environment, but also show that it is not impervious to full-scale algorithmic collusion. These findings contribute to the ongoing debates about algorithmic pricing and its implications.
Igor Sadoune1,2,∗, Marcelin Joanis1,2,‡, Andrea Lodi3,‡
1Department of Mathematics and Industrial Engineering, Polytechnique Montreal, Montreal, QC, Canada
2CIRANO, Montreal, QC, Canada
3Jacobs Technion-Cornell Institute, Cornell Tech and Technion - IIT, New York, NY, USA
Corresponding author: igor.sadoune@polymtl.ca.
These authors contributed equally to this work: marcelin.joanis@polymtl.ca, andrea.lodi@cornell.edu.
@@ 1 Introduction
Concerns about algorithmic pricing have grown as technological advancements in machine learning, widespread data availability, and the digitization of the economy have created fertile grounds for such practices [1, 2, 3, 4]. Algorithmic pricing can potentially lead to algorithmic collusion, as observed in digital markets such as online retail [5, 6] and electricity supply markets [7, 8, 9, 10, 11]. These sectors demonstrate the potential for artificial decision-makers to learn collusive behaviors through dynamic pricing algorithms and optimal bidding strategies. This echoes the concerns about algorithmic collusion in public procurement, which are increasingly relevant, especially as digital platforms and automated decision-making tools become more prevalent in these settings. Since public procurement is a crucial mechanism for collective welfare, the vulnerability of this sector to collusive outcomes is particularly alarming. For example, it has been shown that bidding patterns were incompatible with a competitive equilibrium in Ukrainian E-procurement data [12]. Given these developments, questions arise. Is public procurement directly threatened by algorithmic pricing? And if so, under which conditions?
@@
In this paper, we introduce the Minimum Price Markov Game (MPMG), a model designed to provide a framework for studying emergent behavior in the context of minimum price auctions, which are commonly used in real-world public procurement. We draw on Robert Axelrod’s concept of "the emergence of cooperation among egoists" [22], exploring tacit coordination among isolated agents in the MPMG. Departing from the traditional game-theoretic Bayesian framework of first-price auctions, we consider a game of complete information based on the assumption of data availability. Indeed, public auction data typically includes the features of auctions and firms for past and ongoing contracts, as well as the associated bids [23]. This allows us to base the MPMG on the Prisoner’s Dilemma (PD), leveraging a structure that captures the essence of public procurement as a social dilemma—competitive by law but cooperative by nature. Indeed, the PD is well known for encapsulating the paradox of rationality central to our discussion on the study of emergent coordination among isolated agents in public markets. Furthermore, we test the MPMG using various state-of-the-art Multi-Agent Reinforcement Learning (MARL) methods and thereby show that the minimum price rule, in our setting, is generally difficult to achieve but far from impervious to tacit coordination among isolated decision-makers.
@@
Our methodological choice, which relies on concepts defined in Section 2, is motivated by extensive literature that shows the effectiveness of combining game theory and reinforcement learning for modeling strategic interactions and optimizing behavior through dynamic adaptation among agents. We explore such literature in Section 3 and show that our methodological framework is ideal for exploring emergent behaviors and potential tacit collusion in public procurement markets. In Section 4, we provide the mathematical model for the MPMG and discuss the benefits of the Markov property in our setting. Section 5 presents the results of our computer experiments and concludes on their implications for tacit coordination in minimum price-driven auctions. Finally, Section 6 provides perspectives for future research and practical considerations.
2 Preliminaries
A Markov Decision Process (MDP) is defined by a tuple $(S,A,P,R,\gamma)$, where $S$ represents a finite set of states and $A$ is a finite set of actions. The state transition probability function $P(s^{\prime}|s,a)$ denotes the probability of moving from state $s$ to state $s^{\prime}$ given action $a$, while the reward function $R(s,a)$ specifies the immediate reward received after performing action $a$ in state $s$. The discount factor $\gamma$ determines the present value of future rewards, balancing immediate and long-term gains. A higher discount factor places greater emphasis on future rewards, encouraging the agent to consider long-term benefits, while a lower discount factor prioritizes immediate rewards. The goal in an MDP is to identify a policy $\pi$ that maximizes the expected cumulative reward. The key feature of an MDP is the Markov property, which asserts that future states depend only on the current state and the actions taken by the agents, not on the sequence of events that preceded it, ensuring that past states are irrelevant once the current state is known.
A Markov game is a generalization of a Markov Decision Process (MDP) to multiple interacting agents. The Markov property extends to this multi-agent setting, distinguishing Markov games from iterated games, which involve repeated play of a single-stage game with a static state representation based on past plays. Formally, a Markov game for $n$ players can be defined by
  • A set of states $\mathcal{S}$
  • For each player $i$, a set of actions $\mathcal{A}_{i}$
  • A transition function $T(s,a_{1},\dots,a_{n},s^{\prime})=\Pr(s^{\prime}\mid s,a_{1},\dots,a_{n})$ specifying the probability of moving from state $s$ to state $s^{\prime}$ given the actions $(a_{1},\dots,a_{n})$ taken by the players
  • For each player $i$, a reward function $R_{i}(s,a_{1},\dots,a_{n})$ specifying the immediate reward received by player $i$ after actions $(a_{1},\dots,a_{n})$ are taken in state $s$
Players aim to maximize their own total expected return, typically discounted over time. The solution concepts for Markov Games often involve finding Nash Equilibria, where no player can improve their payoff by unilaterally changing their strategy.
Multi-Agent Reinforcement Learning (MARL) extends single-agent reinforcement learning (RL) techniques to settings with multiple agents learning concurrently in a shared environment. In RL, an agent interacts with its environment to maximize cumulative rewards, where key components include agent, environment, state, action, reward, policy, and value function. The agent, acting as the decision-maker, responds to the environment that consists of external factors and situations represented as states. Actions are possible decisions the agent can take, evaluated by rewards as feedback on their success. The policy is the agent’s strategy for selecting actions based on the current state, while the value function estimates future rewards to determine the desirability of states or actions. In MARL, multiple agents simultaneously learn and update their policies, causing the environment to appear non-stationary from any single agent’s perspective. This dynamic leads to more complex interactions compared to single-agent RL, requiring robust techniques to address the challenges of Markov games and multiagent systems in general.
Multi-Armed Bandit (MAB) algorithms represent a simplified form of RL or MARL, focusing on the exploration-exploitation trade-off without considering future states, that is, without a model of the environment. The concept of "arms" in multi-armed bandits is analogous to the action space available to each agent. Each "arm" or action represents a distinct stochastic process with its own, a priori unknown, probability distribution of rewards.
Q-learning is an off-policy algorithm for RL and MARL problems where agents learn the value of actions directly by using a Q-value function to predict the expected utility of taking a given action in a given state. The Q-value function, denoted as $Q(s,a)$, represents the expected future rewards when an agent takes action $a$ in state $s$ and follows the optimal policy thereafter. The agent updates these values based on the rewards received. Off-policy means that the agent can learn the value of the optimal policy independently of the agent’s current actions, typically by using a replay buffer to store and reuse past experiences. This allows the agent to learn from data generated by different policies or even random actions, not just the current policy it is following. This class of methods is widely used for discrete action space problems because, in contrast to continuous problems, the state-action pair provides a clear and direct signal for learning.
Policy gradient methods focus on optimizing the policy directly by gradient ascent on expected rewards for RL or MARL problems. In MARL, policy gradient techniques, including the actor-critic framework, are used by each agent to adjust their policies in a way that considers the impact of other agents’ strategies, often involving complex inter-agent coordination and the handling of high-dimensional action spaces. The policy $\pi_{\theta}$ is updated by gradient ascent; namely
$\nabla J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)Q^{\pi}(s,a)\right],$
where $\theta$ represents the parameters of the policy, and $Q^{\pi}(s,a)$ is the expected return following action $a$ in state $s$ under policy $\pi$.
@@ 3 Related Work
MDP and auction design. The MDP framework facilitates the exploration, understanding, and prediction of strategic interactions and behaviors within market designs. Examples include computing Nash equilibria in auction games [24, 25] and optimizing bidding strategies in combinatorial auctions [26, 27]. Although the Markov game framework is influential, bandit algorithms have also proven to be exceptionally suitable for modeling optimal bidding strategies in online auction designs [28, 29, 30, 31]. This is due to the simplicity and effectiveness of bandit algorithms, which operate in a stateless or single-state environment and make decisions independently of past actions without the need for an underlying state transition model.
@@
MDP and Stackelberg game. Stackelberg models, in which players move sequentially rather than simultaneously and therefore were historically used to study hierarchical or sequential leadership, are being increasingly used in various applications within the Markov game and multiagent learning framework [32]. The Stackelberg game is a staple in the analysis of various economic and strategic situations where the ability to move first or commit to a decision before one’s competitors can confer a strategic advantage. This includes industries with clear market leaders who can influence the market environment [33]. For example, RL agents have been trained to learn the Stackelberg equilibrium in a Markov decision process representing economic mechanisms for assigning items to individualistic agents in sequential-pricing stages [34, 35, 36].
@@
Multiagent learning, social dilemmas, and tacit coordination. In the exploration of tacit collusion and cooperative strategies within repeated game frameworks, a multitude of studies have contributed to a nuanced understanding of how agents can implicitly coordinate their actions under restricted communication conditions to optimize collective outcomes. The effectiveness of collusion in auction settings without explicit communication is a recurring theme [37], with similar phenomena observed in oligopoly models through Q-learning mechanisms where collusion emerges without direct interaction among firms [38, 39]. These findings are complemented by insights into multi-agent control using deep reinforcement learning, which further illustrates how complex cooperative behaviors can be engineered even in large-scale agent systems [40].
@@
In scenarios where direct communication is either limited or non-existent, learning algorithms play a pivotal role in facilitating implicit cooperation and understanding competitive behaviors. For instance, learning methods that incorporate the anticipation of other agents’ learning processes have shown promise in fostering cooperative strategies that are robust against exploitation [41]. This is particularly evident in the Iterated Prisoner’s Dilemma, where strategies developed through evolutionary algorithms have been demonstrated to outperform traditional approaches [42]. Indeed, the contradictory nature of these games makes their study with RL agents quite challenging, as demonstrated in [43]. Moreover, the development of frameworks that enforce cooperation and resist collusion highlights the potential for creating stable cooperative environments in public goods games and other multi-agent settings [44, 45]. These strategies ensure that cooperation is not only a temporary state but a stable equilibrium, which is crucial for the long-term sustainability of collective actions in decentralized systems.
@@ 4 The Minimum Price Markov Game
In public procurement, there is a tension between the pursuit of personal gain and collective welfare. Firms strive to maximize their individual profits, while the state aims to allocate public funds efficiently and maintain market competition and fairness. From the firms’ perspective, individually rational choices made without cooperation can lead to suboptimal outcomes for all participants, as maximal profits are only achievable through cooperative behavior. This conflict between individual and collective rationality is the core characteristic of social dilemmas. The minimum price rule, with its mathematical elegance, does not diminish this inherent conflict; rather, it preserves this tension fully, maintaining the balance between self-interest and group [^1] For clarity, here, the term ”group” refers to the group of firms, in contrast to the terms ”common” or ”collective”, which we employ to denote the whole economy. benefit.
It follows that a minimum price-ruled auction game formulation will fall in accordance with the principles of social dilemma games. Namely, it must display non-zero-sum interactions for which players can benefit from cooperating or suffer from a single defection. The dominant strategy (Nash equilibrium), being collectively irrational, is thus Pareto inefficient, meaning that the group optimal outcome aligns with a Pareto equilibrium. An efficient coordination scheme leads to the strategy profile associated with the profit-maximizing Pareto equilibrium. In other words, the role of coordination is to transition from the Nash to a Pareto equilibrium.
In this section, we provide a formulation for the MPMG, which we believe is a reasonable approximation of the typical real-world minimum price-ruled public procurement process. Our formulation is grounded in assumptions that capture the core essence of this process, creating a conducive environment to observe how artificial learners dynamically shape their strategies in such dilemma. The MPMG is founded on a static normal form game, the Minimum Price Game (MPG), which we will define first.
@@ 4.1 The Minimum Price Game: A Single-Stage Formulation
The MPG involves a public contractor, implicitly represented by the game’s rules and payoff structure, along with $n$ firms competing to win a contract in a minimum price auction, where the lowest bidder wins. Let $\mathcal{N}=\{1,\dots,n\}$ represent the set of firms in a specific industry. Each firm $i$ has a market power defined by the parameter $\beta_{i}\in(0,1)$, with $\sum_{i\in\mathcal{N}}\beta_{i}=1$, indicating that the average market strength, $\mu(\beta)$, is always $\frac{1}{n}$. This market power might represent the size of the firm or its market share, influencing its ability to reduce costs and leverage economies of scale. All firms share the same cost function, which determines their bids
$b_{i}=(1-\beta_{i})v\quad\forall i\in\mathcal{N},$ (1)
where $v$ is the value of the contract. The notion of strong and weak firms is relative to the distribution of $\beta$, but according to Equation (1), the closer $\beta$ is to 1, the lower the firm can bid, thus the stronger it is. Market heterogeneity is quantified by $\sigma(\beta)$, the standard deviation of the distribution of power parameter values. In a perfectly homogeneous market, $\sigma(\beta)=0$ and $\beta_{i}=\frac{1}{n}$ for all $i$. In such cases, the set of bids could display multiple minima if firms submit their prices directly following Equation (1). In other words, the homogeneous case is analogous to a game with no private values. This possibility arises from strategic considerations, though we have not yet fully outlined the space of strategies. We denote by $\Omega\subseteq\mathcal{N}$ the set of firms associated with the lowest bids.
Assumption 1 (Single Public Contractor).
All public contractors share the same goal.
In real-world scenarios, the set of public contractors in a given procurement market usually includes many entities. Despite this plurality, it is safe to assume that all public contractors work towards the same overarching goal. Assumption 1 allows us to reduce the set of public contractors to a singleton, a realistic simplification that will help in formulating an elegant model.
@@
Assumption 2 indicates that payoffs are determined in a single round of bid submissions. This necessitates that the payoff function accommodates the possibility of $\Omega$ including multiple firms. A simple solution for handling ties is to randomly select a winner, either by using the uniform distribution ($i^{*}\sim\mathcal{U}(\Omega)$) or by adding a stochastic perturbation to the bids ($b_{i}+\epsilon_{i}$). However, random selection would lead to a more complicated stochastic game without realistic justification. Furthermore, the process of random selection in ties forces $\Omega$ into a singleton, which may be overly restrictive. In practice, the likelihood of a tie in such a scenario is relatively low, and when it occurs, discrimination is not random.
Instead, we advocate for shared payoffs in the event of a tie. Sharing payoffs among the lowest bidders is justified for two reasons. First, even from a static perspective, this approach preserves and highlights the tension between individual and group rationality, which is the feature essential to our formulation and endeavor. Second, the MPG intends to provide a theoretical basis for further dynamic implementations. In dynamic settings, distributing the payoff among several bidders in a tie situation is analogous to random winner selection, as the cumulative payoffs for each bidder equalize over multiple auctions for a given repeated strategy profile.
@@
All firms submit their bids simultaneously.
Although bids are, in practice, collected over a certain period, they remain sealed and confidential. This ensures that all bids are (virtually) submitted at the same time, thus justifying Assumption 3.
@@
The contract value is common for all firms.
@@
Assumption 4 simplifies the commonly used Bayesian game formulation of first-price auction to reflect the conditions of Bertrand competition [46], where firms compete on prices in a transparent market. This assumption is supported by the uniformity of contract requirements and regulatory costs across bidders, market transparency, and the homogeneity in the evaluation of contract values. These factors standardize the perceived value of the contract, thereby aligning all firms’ bids based on a common economic assessment rather than individual speculative valuations. Assumptions 1 to 4, combined with the simple one-dimensional minimum price rule, lead to a deterministic normal-form game formulation.

### v2

@@ Abstract
This paper introduces the Minimum Price Markov Game (MPMG), a theoretical model that reasonably approximates real-world first-price markets following the minimum price rule, such as public auctions. The goal is to provide researchers and practitioners with a framework to study market fairness and regulation in both digitized and non-digitized public procurement processes, amid growing concerns about algorithmic collusion in online markets. Using multi-agent reinforcement learning-driven artificial agents, we demonstrate that (i) the MPMG is a reliable model for first-price market dynamics, (ii) the minimum price rule is generally resilient to non-engineered tacit coordination among rational actors, and (iii) when tacit coordination occurs, it relies heavily on self-reinforcing trends. These findings contribute to the ongoing debate about algorithmic pricing and its implications.
Keywords Algorithmic Game Theory  $\cdot$ Multiagent Reinforcement Learning  $\cdot$ Algorithmic Coordination  $\cdot$ Algorithmic Pricing
To replicate the study, follow GitHub.
@@ 1 Introduction
Concerns about algorithmic pricing have grown as technological advancements in machine learning, widespread data availability, and the digitization of the economy have created fertile grounds for such practices [1, 2, 3, 4]. Algorithmic pricing can potentially lead to collusion, as observed in digital markets such as online retail [5, 6] and electricity supply markets [7, 8, 9, 10, 11]. These sectors demonstrate the potential for artificial decision-makers to learn collusive behaviors through dynamic pricing algorithms and optimal bidding strategies. This echoes the concerns about algorithmic collusion in public procurement, which are increasingly relevant, especially as digital platforms and automated decision-making tools become more prevalent in these settings. Since public procurement is a crucial mechanism for collective welfare, the vulnerability of this sector to collusive outcomes is particularly alarming. For example, it has been shown that bidding patterns were incompatible with a competitive equilibrium in Ukrainian E-procurement data [12]. Given these developments, questions arise. Is public procurement directly threatened by algorithmic pricing? And if so, under which conditions?
@@
In this paper, we introduce the Minimum Price Markov Game (MPMG), a model designed to provide a framework for studying emergent behavior in the context of minimum price auctions, which are commonly used in real-world public procurement. We draw on Robert Axelrod’s concept of “the emergence of cooperation among egoist” [22], exploring tacit coordination among isolated agents in the MPMG. Departing from the traditional game-theoretic Bayesian framework of first-price auctions, we consider a game of complete information based on the assumption of data availability. Indeed, public auction data typically includes the features of auctions and agents from past to ongoing contracts, as well as the associated bids [23]. This allows us to base the MPMG on the Prisoner’s Dilemma (PD), leveraging a structure that captures the essence of public procurement as a social dilemma—competitive by law but cooperative by nature. Indeed, the PD is well known for encapsulating the paradox of rationality central to our discussion on the study of emergent coordination among isolated agents in public markets. Furthermore, we test the MPMG using various state-of-the-art Multi-Agent Reinforcement Learning (MARL) methods and thereby show that the minimum price rule, in our setting, is generally difficult to achieve but far from impervious to tacit coordination among isolated decision-makers.
@@
Our methodological choice is motivated by extensive literature that shows the effectiveness of combining game theory and reinforcement learning for modeling strategic interactions and optimizing behavior through dynamic adaptation among agents. We explore such literature in Section 2 and show that our methodological framework is ideal for exploring emergent behaviors and potential tacit collusion in public procurement markets. In Section 3, we provide the mathematical model for the MPMG and discuss the benefits of the Markov property in our setting. Section 4 presents the results of our computer experiments and concludes on their implications for tacit coordination in minimum price-driven auctions. Finally, Section 5 provides interpretation, practical considerations and perspectives.
@@ 2 Related Work
The MDP framework facilitates the exploration, understanding, and prediction of strategic interactions and behaviors within market designs. Examples include computing Nash equilibria in auction games [24, 25] and optimizing bidding strategies in combinatorial auctions [26, 27]. Although the Markov game framework is influential, bandit algorithms have also proven to be exceptionally suitable for modeling optimal bidding strategies in online auction designs [28, 29, 30, 31]. This is due to the simplicity and effectiveness of bandit algorithms, which operate in a stateless or single-state environment and make decisions independently of past actions without the need for an underlying state transition model.
MDP and Stackelberg game.
@@
Stackelberg models, in which players move sequentially rather than simultaneously and therefore were historically used to study hierarchical or sequential leadership, are being increasingly used in various applications within the Markov game and multiagent learning framework [32]. The Stackelberg game is a staple in the analysis of various economic and strategic situations where the ability to move first or commit to a decision before one’s competitors can confer a strategic advantage. This includes industries with clear market leaders who can influence the market environment [33]. For example, RL agents have been trained to learn the Stackelberg equilibrium in a Markov decision process representing economic mechanisms for assigning items to individualistic agents in sequential-pricing stages [34, 35, 36].
Multiagent learning, social dilemmas, and tacit coordination.
@@
In the exploration of tacit collusion and cooperative strategies within repeated game frameworks, a multitude of studies have contributed to a nuanced understanding of how agents can implicitly coordinate their actions under restricted communication conditions to optimize collective outcomes. The effectiveness of collusion in auction settings without explicit communication is a recurring theme [37], with similar phenomena observed in oligopoly models through Q-learning mechanisms where collusion emerges without direct interaction among agents [38, 39]. These findings are complemented by insights into multi-agent control using deep reinforcement learning, which further illustrates how complex cooperative behaviors can be engineered even in large-scale agent systems [40].
@@
In scenarios where direct communication is either limited or non-existent, learning algorithms play a pivotal role in facilitating implicit cooperation and understanding competitive behaviors. For instance, learning methods that incorporate the anticipation of other agents’ learning processes have shown promise in fostering cooperative strategies that are robust against exploitation [41]. This is particularly evident in the Iterated Prisoner’s Dilemma, where strategies developed through evolutionary algorithms have been demonstrated to outperform traditional approaches [42]. Indeed, the contradictory nature of these games makes their study with RL agents quite challenging, as demonstrated in [43]. Moreover, the development of frameworks that enforce cooperation and resist collusion highlights the potential for creating stable cooperative environments in public goods games and other multi-agent settings [44, 45].
@@ 3 The Minimum Price Markov Game
In real-world multiagent systems, a common observation is that heterogeneous agents coexist in non-cooperative environments. For example, in heterogeneous markets, stronger agents typically enjoy a competitive advantage, making growth, or even survival, more difficult for weaker ones. This disparity in market power—whether due to factors like market share, technology, or cost structures—creates a stratification of prices, influencing price equilibrium. In the context of first-price public auctions, stronger agents tend to submit the lowest bids, leading to the expectation that they will win more frequently. This assumption is reasonable, as their ability to cut costs is positively correlated with their operational capacity.
In a game-theoretic setting, it is necessary to define a payoff function that applies to both heterogeneous and homogeneous markets without losing generality. The payoff structure should smoothly transition from homogeneous to heterogeneous markets, allowing for a mathematically consistent analysis of strategic behaviors across different numbers of players. Market heterogeneity should be reflected in differences in payoffs, win frequencies, and incentives to avoid price collusion, with these variations proportionate to the gaps in market strength.
Furthermore, in public procurement, there is a tension between the pursuit of personal gain and collective welfare. Firms strive to maximize their individual profits, while the state aims to allocate public funds efficiently and maintain market competition and fairness. From the agents’ perspective, individually rational choices made without cooperation can lead to suboptimal outcomes for all participants, as maximal profits are only achievable through cooperative behavior. This conflict between individual and collective rationality is the core characteristic of social dilemmas. The minimum price rule, with its mathematical elegance, does not diminish this inherent conflict; rather, it preserves this tension fully, maintaining the balance between self-interest and group benefit.
It follows that a minimum price-ruled auction game formulation will fall in accordance with the principles of social dilemma games. Namely, it must display non-zero-sum interactions for which players can benefit from cooperating or suffer from a single defection. The dominant strategy (Nash Equilibrium), being collectively irrational, is thus Pareto inefficient, meaning that an efficient coordination scheme leads to the strategy profile associated with the non-Nash Pareto Optimal outcome. In other words, the role of coordination is to transition from the Nash Equilibrium to the profit-maximizing strategy profile.
In addition, the single-stage formulation must capture the core elements of market dynamics, which inherently unfold through repeated interactions. These characteristics must be maintained in its dynamic extension, such as in the iterated MPG and the Minimum Price Markov Game (MPMG).
In this section, we provide a formulation for the MPMG, which we believe is a reasonable approximation of the typical real-world minimum price-ruled public procurement process. Our formulation is grounded in assumptions that capture the core essence of this process, creating a conducive environment to observe how artificial learners dynamically shape their strategies in such dilemma. The MPMG is founded on a static normal-form game, the Minimum Price Game (MPG), which we will define first.
@@ 3.1 The Minimum Price Game: A Single-Stage Formulation
The MPG involves $n$ agents, represented by a set $\mathcal{N}=\{1,\dots,n\}$, competing to win a contract valued at $v$ in a first-price procurement auction.
@@
All agents submit their bids simultaneously.
Although bids are, in practice, collected over a certain period, they remain sealed and confidential. This ensures that all bids are (virtually) submitted at the same time, thus justifying Assumption 2. Assumptions 1 and 2 complement each other and allow for a normal-form game formulation.
@@
The contract value is common for all agents.
@@
Assumption 3 simplifies the commonly used Bayesian game formulation of first-price auction to reflect the conditions of Bertrand competition [46], where agents compete on prices in a transparent market. This assumption is supported by the uniformity of contract requirements and regulatory costs across bidders, market transparency, and the homogeneity in the evaluation of contract values. These factors standardize the perceived value of the contract, thereby aligning all agents’ bids based on a common economic assessment rather than individual speculative valuations.
Strategies.